OpenStudy (anonymous):

The length of a rectangle is 6 m, and its width is 14 m. What is the area of half of the rectangle? 84 m2 102 m2 42 m2 22 m2

OpenStudy (anonymous):

pleeeeeeaaasse help

OpenStudy (anonymous):

Full Area of Rectangle is given by: \[\large Area = Length \times Width\] So for half divide it by 2: \[\large \color{green}{Area = \frac{1}{2} \times Length \times Width}\]

OpenStudy (anonymous):

thanks so much! (: 42.

OpenStudy (anonymous):

So here Length is 6 and Width = 14 So plug in the values and tell me what do you get??

OpenStudy (anonymous):

Well Done.. 42 is correct..